# IsoSEL: Isometric Structural Entropy Learning for Deep Graph Clustering in Hyperbolic Space

### v1

@@ VIII-G Embedding Expressiveness of IsoSEL
(j) IsoSEL:prediction
Figure 5: Visualization of real and predicted data.
@@ VIII-H Case Study on Imbalanced Clusters
We examine the clustering boundaries among imbalanced clusters. Real graphs often exhibit skewed class distributions, and for this case study, we use the Photo dataset. The proportion of $8$ clusters in Photo is $25\%$, $22\%$, $11\%$, $11\%$, $10\%$, $9\%$, $4\%$ and $4\%$. We compare IsoSEL with gCool, DGCluster, RGC and LSEnet, and visualize the node embedding in Fig. 5 where different clusters are denoted by different colors. The distributions of true labels and predicted labels are given in the top and bottom subfigures, respectively. We invoke the normal t-sne for gCool, DGCluster, RGC, while employing the hyperbolic metric for LSEnet and IsoSEL, so that all the visualizations are finally presented in 2D Euclidean space. In Fig. 5, we obtain three findings as follows. (i) Clustering model without $K$ may fail to uncover the correct cluster number, e.g., RGC outputs $7$ node clusters on Photo dataset, but IsoSEL gives the true cluster number $8$. (ii) We observe that deep clustering models tend to generate biased clustering boundary. For example, in RGC, the members of violet cluster and the minority gray cluster (at the upper left in Fig. 5c) are assimilated into the majority cluster colored in purple, comparing the true label in Fig. 5h. The reason lies in that the deep models typically perform clustering in the representation space and minority clusters are underrepresented in the self-supervised learning owing to limited samples. (iii) Different from other deep models, the minority clusters are easier to be identified and discovered in both LSEnet and IsoSEL. We highlight that, the differentiable structural information objective effectively discovers node clusters without $K$ and is suitable to identify minority clusters in imbalanced graphs. In addition, compared to LSEnet, IsoSEL further mitigates the bias in clustering boundary with the hyperbolic reweighing proposed in Lorentz tree contrastive learning.
@@ VIII-I Visualizing Hyperbolic Partitioning Tree
We visualize the hyperbolic partitioning trees of IsoSEL. In IsoSEL, the partitioning tree is created in the Lorentz model of hyperbolic space. Note the fact that the 2D Lorentz model is represented in 3D Euclidean space and is hard to be visualized. To bridge this gap, we perform the stereographic projection $\Psi$ [55] on the Lorentz model and obtain the equivalent Poincaré disc $\mathbb{B}^{2}$, a 2D open disc that is more preferable for visualization. Given the standard curvature $\kappa=-1$, for any point in the Lorentz model $\boldsymbol{x}=[x_{1}\ \boldsymbol{x}_{s}]^{\top}\in\mathbb{L}^{2}$, we derive the corresponding point in Poincaré disc via $\Psi(\boldsymbol{x})=\frac{\boldsymbol{x}_{s}}{x_{1}+1}\in\mathbb{B}^{2}$. With $\Psi$, we study the clustering results on two toy examples of Karate and Football datasets, and show the visualization in the 2D Poincaré disc in Fig. 6. The partitioning tree with learnt clusters are plotted in Fig. 6 (a) and (b), while true labels are given in Fig. 6 (c) and (d). Two important findings are summarized as follows: (i) The clustering results of IsoSEL are generally aligned with the ground truth distribution. (ii) The learnt embeddings of each clusters are well separated, showing the effectiveness of IsoSEL.
@@ X Conclusion
Our work bridges structural information with deep learning and provides a principled solution for graph clustering without without requiring a predefined number of clusters while accounting for the imbalance in real-world graphs. Theoretically, we establish differentiable structural information in continuous realm, and prove its effectiveness in clustering without $K$ and detecting minority clusters. Accordingly, we present the IsoSEL framework where LSEnet is designed to learn a deep partitioning tree via gradient backpropagation in the Lorentz model of hyperbolic space. Furthermore, we propose Lorentz tree contrastive learning with isometric augmentation to refine the deep tree, thereby enhancing cluster assignment. Extensive experiments demonstrate that IsoSEL effectively generates node clusters and mitigates the bias in the clustering boundary among imbalanced clusters.

### v2

@@ VIII-G Embedding Expressiveness of ASIL
(j) ASIL:prediction
Figure 4: Visualization of real and predicted data on Photo dataset.
@@ VIII-H Case Study on Imbalanced Clusters
We examine the clustering boundaries among imbalanced clusters. Real graphs often exhibit skewed class distributions, and for this case study, we use the Photo dataset. The proportion of $8$ clusters in Photo is $5\%, 22$$2\%$, $\%, 1111\%$, $10\%$, $9\%$, $4\%$ and $4\%$. We compare ASIL with gCool, DGCluster, RGC and LSEnet, and visualize the node embedding in Fig. 4 where different clusters are denoted by different colors. The distributions of true labels and predicted labels are given in the top and bottom subfigures, respectively. We invoke the normal t-sne for gCool, DGCluster, RGC, while employing the hyperbolic metric for LSEnet and ASIL, so that all the visualizations are finally presented in 2D Euclidean space. In Fig. 4, we obtain three findings as follows. (i) Clustering model without $K$ may fail to uncover the correct cluster number, e.g., RGC outputs $7$ node clusters on Photo dataset, but ASIL gives the true cluster number $8$. (ii) We observe that deep clustering models tend to generate biased clustering boundary. For example, in RGC, the members of violet cluster and the minority gray cluster (at the upper left in Fig. 4c) are assimilated into the majority cluster colored in green, comparing the true label in Fig. 4h. The reason lies in that the deep models typically perform clustering in the representation space and minority clusters are underrepresented in the self-supervised learning owing to limited samples. (iii) Different from other deep models, the minority clusters are easier to be identified and discovered in both LSEnet and ASIL. We highlight that, the differentiable structural information objective effectively discovers node clusters without $K$ and is suitable to identify minority clusters in imbalanced graphs. In addition, compared to LSEnet, ASIL further mitigates the bias in clustering boundary with better hyperbolic partitioning tree representations.
@@ VIII-I Visualizing Hyperbolic Partitioning Tree
We visualize the hyperbolic partitioning trees of ASIL. In ASIL, the partitioning tree is created in the Lorentz model of hyperbolic space. Note the fact that the 2D Lorentz model is represented in 3D Euclidean space and is hard to be visualized. To bridge this gap, we perform the stereographic projection $\Psi$ [2] on the Lorentz model and obtain the equivalent Poincaré disc $\mathbb{B}^{2}$, a 2D open disc that is more preferable for visualization. Given the standard curvature $\kappa=-1$, for any point in the Lorentz model $\boldsymbol{x}=[x_{1}\ \boldsymbol{x}_{s}]^{\top}\in\mathbb{L}^{2}$, we derive the corresponding point in Poincaré disc via $\Psi(\boldsymbol{x})=\frac{\boldsymbol{x}_{s}}{x_{1}+1}\in\mathbb{B}^{2}$. With $\Psi$, we study the clustering results on two toy examples of Karate and Football datasets, and show the visualization in the 2D Poincaré disc in Fig. 5. The partitioning tree with learnt clusters are plotted in Fig. 5 (a) and (b), while true labels are given in Fig. 5 (c) and (d). Two important findings are summarized as follows: (i) The clustering results of ASIL are generally aligned with the ground truth distribution. (ii) The learnt embeddings of each clusters are well separated, showing the effectiveness of ASIL.
@@ X Conclusion
Our work presents a principled solution for graph clustering from an information-theoretic perspective, considering the practical challenges of unknown cluster number and imbalanced node distribution. To this end, we extend the classic structural entropy to the continuous realm, so that its partitioning tree can be learned via gradient backpropagation, uncovering clustering structures. Theoretically, we demonstrate its capacity in debiased clustering and connect it to contrastive learning, allowing for effective and efficient clustering via Augmented Structural Information Learning (ASIL). The key innovation is the seamless integration of partitioning tree construction and contrastive learning in hyperbolic space. Without requiring the number of clusters, ASIL achieves debiased node clustering with linear computational complexity in graph size. Empirically, we validate the superiority of ASIL against $20$ strong baselines on multiple benchmark datasets.